Charles has $2.75 in dimes and nickles. The number of nickles is one less than twice the number of dimes.How many of each type of coin does he have?
Answer by TimothyLamb(4379)

5n + 10d = 275
n = 2d - 1
---
5n + 10d = 275
5(2d - 1) + 10d = 275
10d - 5 + 10d = 275
20d = 280
---
d = 14
n = 27
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php